# A similarity reduction

### 17 Jan 2024 v1

@@ 1 Introduction
The Solovév solution [4] and the Herneger-Maschke solution [5] are among the most widely employed analytical solutions of the GSE, the former corresponding to toroidal current density non-vanishing on the plasma boundary while the later to toroidal current density vanishing thereon. In the presence of flow, which plays a role in the transition to improved confinement modes in tokamaks, the equilibrium satisfies a GGSE in general coupled with a Bernoulli equation involving the pressure [6, 7]. For incompressible flow, the density becomes a surface quantity and we obtain the GGSE [7, 8]. Generalized Solovév solutions of the GGSE were derived in [8, 9] and asymptotic expansion solutions to the general linearized GGSE in [10, 11].
A symmetry group of a partial differential equation is a set of transformations of the independent or/and dependent variables that allows to generate new classes of solutions from a single solution known. Symmetry properties of a system of Euler-type equations were studied in [12]. Symmetry properties and solutions to the GSE and the GGSE were studied in [13, 14], whereas symmetry methods for differential equations are exposed in [15]. Translationally symmetric force-free states are constructed in [16]. The literature on exact solutions to the GSE is extensive and a brief reference list includes [17, 18], while other symmetry and similarity reduction methods were employed in [19, 20, 21, 22].
Aim of the present study is to construct new families of analytic solutions to the GGSE by using the method of similarity reduction along the lines of [1], describing equilibrium configurations relevant to experimental fusion and space plasmas. The structure of the paper is as follows: In section 2 we present the direct method for finding the similarity reduction of the GGSE and show that the generalized Solovév solution is a particular case of the above mentioned set of solutions. In section 3 we extend the set of solutions to the GGSE and show that all group-invariant solutions of [3] are in fact particular cases of the extended set of solutions. In section 4 we present a new similarity reduction of the GGSE, not considered in [1], associated with solutions which can describe part of planetary magnetospheres. Finally, in section 5 we present a brief discussion on the novel results of this paper.
@@ 2 Axisymmetric equilibria with non-parallel flow from similarity reduction
Here, employing cylindrical coordinates($z,r,\phi$), the function $u(r,z)$ relates to the poloidal magnetic flux function and labels the magnetic surfaces; the flux function $X(u)$ relates to the toroidal magnetic field; $P_{s}(u)$ is the pressure in the absence of flow; $\rho(u)$ is the density; $M_{p}(u)$ is the Mach function of the poloidal velocity with respect to the respective Alfvén velocity; and $\Phi(u)$ is the electrostatic potential related to the component of the electric field non parallel to the magnetic field. The flux functions $X(u)$, $M_{p}(u)$, $P_{s}(u)$, $\rho(u)$ and $\Phi(u)$ remain arbitrary.
Choosing the free function terms to be constant, i.e., ($(1/2)d/du[X^{2}(1-M_{p}^{2})]=G$, $dP_{s}/du=F$, $(1/2)d/du[\rho(d\Phi/du)^{2}]=E$ with $G$, $F$, $E$ constants) we have the following form of the GGSE that will be the focus of this study
@@
The respective equilibrium configuration inside the separatrix, up-down symmetric with a pair of X-points, is plotted in Fig. 1.
@@
$\beta(x_{r}^{2}+x_{z}^{2})w^{{}^{\prime\prime}}$ $+$ $\left[\left(2\beta_{r}-\frac{1}{r}\beta\right)x_{r}+2\beta_{z}x_{z}+\beta(x_{rr}+x_{zz})\right]w^{{}^{\prime}}+$ (5)
@@
This equation can be satisfied by choosing $x(z)=z^{2}/2$, provided that the functions $\alpha,\beta$ satisfy the ordinary differential equations
@@
For $\nu=1$ we obtain the solution of Eq. (12) in the following two equivalent forms
@@
$u(r,z)=\alpha(r,z)+r[c_{1}[x+\sqrt{x^{2}+1}]+c_{2}[x+\sqrt{x^{2}+1}]^{-1}]\,,$ (15)
where $\alpha(r,z)$ is any solution of the GGSE such as the generalized Solovév solution and $x=z/r$. It is plotted in Fig. 3, for $E=-2,$ $G=0.2$, $F=-4,$ $c_{1}=c_{2}=0.05$. The bounding flux surface corresponds to $U_{b}=0.40$.
@@
For $\nu=1/2$ and $\nu=3/2$ Eq. (12) can’t be solved analytically and it has been solved numerically. The solutions are plotted in Figs. 4, 5, 6 and 7. Specifically, for $\nu=1/2$ we integrated numerically Eq. (12) in the interval $-5\leq x\leq 5$ for $w(-5)=-1$ and $w^{{}^{\prime}}(-5)=1$. Numerical fitting with a polynomial of ninth order gives
@@
The respective solution of Eq. (4) for $E=-2$, $G=0.2$, $F=-4$, is plotted in Fig. 5, where $u=u_{Sol}+cr^{1/2}w$ with $u_{Sol}$ being the generalized Solovév solution (3) and $c=5\times 10^{-4}$.
Figure 5: Up-down assymetric equilibrium determined by (4) with the aid of the numerical solution of Eq. (12) for $w(x)$ with $\nu=1/2$, shown in Fig. 4.
@@
Figure 7: Up-down assymetric equilibrium determined by (4) with the aid of the numerical solution of Eq. (12) for $w(x)$ with $\nu=3/2$, shown in Fig. 6.
It is noted that the equilibrium configurations in Figs. 5 and 7 are up-down asymmetric.
@@ 3 More equilibria from similarity reduction
For a general $\nu$ Eq. (12) has precisely the general solution [1]
$w(x)$ $=$ $c_{1}(1+x^{2})^{\frac{(2\nu-1)}{4}}P_{1/2}^{\frac{(1-2\nu)}{2}}(ix)+$ (17)
$+$ $c_{2}(1+x^{2})^{\frac{(2\nu-1)}{4}}Q_{1/2}^{\frac{(1-2\nu)}{2}}(ix)\,,$
in terms of the associated Legendre functions $P_{1/2}^{(1-2\nu)/2}(ix)$ and $Q_{1/2}^{(1-2\nu)/2}(ix)$ New non-trivial solutions to the GGSE can be generated from any solution to Eq. (13). For example, we will employ the following solutions to Eq. (13):
@@
Moreover if $u(r,z)$ is a solution to the GGSE (2) then $u(r,z)+\beta(r,z)w[x(r,z)]$ is also a solution, thus making it possible to combine solutions with different $\nu$. As an example, combining the solutions for $\nu=3$ with those for $\nu=1$ presented previously and Eq. (18) we find
@@ 5 Conclusions
We have considered a similarity reduction of the GGSE along the lines presented in [1]. New classes of exact solutions to the GGSE were produced, with the generalized Solovév solution being a particular case, describing a variety of D-shaped equilibrium configurations exhibiting up-down symmetry or asymmetry and either positive or negative triangularity. All group-invariant solutions of the GGSE presented in [3] are shown to be particular cases of the above formalism. In addition, a new class of solutions, not considered in [1], has been constructed which can describe part of a planetary magnetoshphere for short distances from the planet surface.
@@
This work has received funding from the National Fusion Programme of the Hellenic Republic--General Secretariat for Research and Innovation.

### 08 Apr 2024 v2

@@ 1 Introduction
The Solovév solution [4] and the Herrnegger-Maschke solution [5] are among the most widely employed analytical solutions of the GSE, the former corresponding to toroidal current density non-vanishing on the plasma boundary while the later to toroidal current density vanishing thereon. In the presence of flow, which plays a role in the transition to improved confinement modes in tokamaks, the equilibrium satisfies a GGSE, in general coupled with a Bernoulli equation involving the pressure [6, 7]. For incompressible flow, the density becomes a surface quantity and we obtain the GGSE [7, 8]. Generalized Solovév solutions of the GGSE were derived in [8, 9] and asymptotic expansion solutions to the general linearized GGSE in [10, 11]. Also, some analytical solutions were recently constructed for the axisymmetric Hall magnetohydrodynamics GSE with incompressible flows [12].
A symmetry group of a partial differential equation (PDE) is a set of transformations of the independent or/and dependent variables that allows to generate new classes of solutions from a single solution known. According to the classical Lie group method, a similarity reduction of a PDE is to determine a one parameter group admitted by the equation. Then one seeks the group-invariant solutions called self-similar solutions (cf. [13] for a recent description of the method). Symmetry properties of a system of Euler-type equations were studied in [14]. Symmetry properties and solutions to the GSE and the GGSE were studied in [15, 16], whereas symmetry methods for differential equations are exposed in [13]. Translationally symmetric force-free states are constructed in [17]. The literature on exact solutions to the GSE is extensive and a brief reference list includes [18, 19], while other symmetry and similarity reduction methods were employed in [20, 21, 22, 23].
Aim of the present study is to construct new families of analytic solutions to the GGSE by using an alternative method of similarity reduction along the lines of [1], describing equilibrium configurations relevant to experimental fusion and space plasmas. Accordingly, a PDE can be reduced to an ordinary differential equation by employing an appropriate functional relation involving the dependent and independent variables, as Eq. (4) below, after imposing certain conditions.
The structure of the paper is as follows: In section 2 we present the direct method for finding the similarity reduction of the GGSE and show that the generalized Solovév solution is a particular case of the above mentioned set of solutions. In section 3 we extend the set of solutions to the GGSE and show that all group-invariant solutions of [3] are in fact particular cases of the extended set of solutions. In section 5 we present a new similarity reduction of the GGSE, not considered in [1], associated with solutions which can describe part of planetary magnetospheres. Finally, in section 5 we present a brief discussion on the novel results of this paper.
@@ 2 Axisymmetric equilibria with non-parallel flow from similarity reduction
Here, employing cylindrical coordinates($z,r,\phi$), the function $u(r,z)$ relates to the poloidal magnetic flux function and labels the magnetic surfaces; the flux function $X(u)$ relates to the toroidal magnetic field (cf. Eq. (3) below); $P_{s}(u)$ is the pressure in the absence of flow; $\rho(u)$ is the density; $M_{p}(u)$ is the Mach function of the poloidal velocity with respect to the respective Alfvén velocity; and $\Phi(u)$ is the electrostatic potential related to the electric field. The flux functions $X(u)$, $M_{p}(u)$, $P_{s}(u)$, $\rho(u)$ and $\Phi(u)$ remain arbitrary. Also, the magnetic field and the velocity are given by the relations
${\mathbf{B}}=I\nabla\phi+\left(1-M_{p}^{2}\right)^{-1/2}\nabla\phi\times\nabla u,\ \ {\mathbf{v}}=\frac{M_{p}}{\sqrt{\rho}}{\mathbf{B}}-r^{2}\left(1-M_{p}\right)^{1/2}\Phi^{\prime}\nabla\phi$ (2)
where
$I=\frac{X}{\left(1-M_{p}\right)}-r^{2}\frac{\sqrt{\rho}M_{p}\Phi^{\prime}}{\left(1-M_{p}\right)^{1/2}}$ (3)
Compared with the GSE the GGSE has the additional $R^{4}$-term related to the non parallel to the magnetic field component of the velocity through the electrostatic potential $\Phi$. This term can significantly modify the equilibrium; for example, in the presence of this term, the well known Solovév equilibrium [4], which is similar to the configuration of Fig. 1, acquires an addition X-point outside the separatrix (cf. Fig. 11 of [8]). Also, although plasma rotation in magnetic confinement systems is low, i.e. $M_{p}\sim 10^{-2}$, the respective electric field and its shear play an important role in the transitions to improved confinement regimes, as the L-H-transition [24].
Choosing the flux function terms to be constant, i.e., ($(1/2)d/du[X^{2}(1-M_{p}^{2})]=G$, $dP_{s}/du=F$, $(1/2)d/du[\rho(d\Phi/du)^{2}]=E$ with $G$, $F$, $E$ constants) we have the following form of the GGSE that will be the focus of this study
@@
The respective equilibrium configuration inside the separatrix, up-down symmetric with a pair of X-points, is plotted in Fig. 1. For $E=0$, (5) reduces to the well known Solovév solution [4].
@@
$\beta\left[(x_{r})^{2}+(x_{z})^{2}\right]w^{{}^{\prime\prime}}$ $+$ $\left[\left(2\beta_{r}-\frac{1}{r}\beta\right)x_{r}+2\beta_{z}x_{z}+\beta(x_{rr}+x_{zz})\right]w^{{}^{\prime}}+$ (7)
@@
This equation can be satisfied by choosing $x(z)=z^{2}/2$, provided that the functions $\alpha,\beta$ satisfy the ODEs
@@
Eq. (14) has the analytic solution [1]
$w(x)$ $=$ $c_{1}(1+x^{2})^{\frac{(2\nu-1)}{4}}P_{1/2}^{\frac{(1-2\nu)}{2}}(ix)+$ (16)
$+$ $c_{2}(1+x^{2})^{\frac{(2\nu-1)}{4}}Q_{1/2}^{\frac{(1-2\nu)}{2}}(ix)\,,$
in terms of the associated Legendre functions $P_{1/2}^{(1-2\nu)/2}(ix)$ and $Q_{1/2}^{(1-2\nu)/2}(ix)$. These functions involve infinite converging series; only for particular values of $\nu$, e.g. $\nu=0,1,2$ to be considered below, the solution is expressed in closed form, i.e. in terms of Legendre polynomials. For this reason, in the former case, instead of employing the analytic solution we preferred to solve Eq. (14) numerically.
In particular, for $\nu=1$ we obtain the solution of Eq. (refeq13) in the following two equivalent forms
@@
$u(r,z)=u_{g}(r,z)+r[c_{1}[x+\sqrt{x^{2}+1}]+c_{2}[x+\sqrt{x^{2}+1}]^{-1}]\,,$ (18)
where $u_{g}(r,z)$ is any solution of the GGSE such as the generalized Solovév solution (5) and $x=z/r$. It is plotted in Fig. 3, for $E=-2,$ $G=0.2$, $F=-4,$ $c_{1}=c_{2}=0.05$. The bounding flux surface corresponds to $U_{b}=0.40$.
@@
For $\nu=1/2$ and $\nu=3/2$ the solution of Eq. (14) is expressed in terms of infinite series; thus, as we already noted earlier, we preferred solving Eq. (14) numerically. The solutions are plotted in Figs. 4, 5, 6 and 7. Specifically, for $\nu=1/2$ we integrated numerically Eq. (14) in the interval $-5\leq x\leq 5$ for $w(-5)=-1$ and $w^{{}^{\prime}}(-5)=1$. Numerical fitting with a polynomial of ninth order gives
@@
The respective solution of Eq. (6) for $E=-2$, $G=0.2$, $F=-4$, is plotted in Fig. 5, where $u=u_{Sol}+cr^{1/2}w$ with $u_{Sol}$ being the generalized Solovév solution (5) and $c=5\times 10^{-4}$. The configuration is D-shaped up-down asymmetric.
Figure 5: Up-down asymmetric equilibrium determined by (6) with the aid of the numerical solution of Eq. (14) for $w(x)$ with $\nu=1/2$, shown in Fig. 4.
@@
Figure 7: Up-down asymmetric equilibrium with a lower X-point determined by (6) with the aid of the numerical solution of Eq. (14) for $w(x)$ with $\nu=3/2$, shown in Fig. 6.
This is a diverted equilibrium with a lower X-point.
@@ 3 More equilibria from similarity reduction
New non-trivial solutions to the GGSE can be generated from any solution to Eq. (15). For example, we will employ the following solutions to Eq. (15):
@@
Moreover if $u_{g}(r,z)$ is a solution to the GGSE (4) then $u_{g}(r,z)+\beta(r,z)w[x(r,z)]$ is also a solution, thus making it possible to combine solutions with different $\nu$. As an example, combining the solutions for $\nu=3$ with those for $\nu=1$ presented previously and Eq. (20) we find
@@ 5 Conclusions
We have considered a similarity reduction of the GGSE along the lines presented in [1]. New classes of exact solutions to the GGSE were produced, with the generalized Solovév solution being a particular case, describing a variety of D-shaped equilibrium configurations exhibiting up-down symmetry or asymmetry and either positive or negative triangularity and diverted configurations with a lower X-point. All group-invariant solutions of the GGSE presented in [3] are shown to be particular cases of the above formalism. In addition, a new class of solutions, not considered in [1], has been constructed which can describe part of a planetary magnetoshphere for short distances from the planet surface.
@@
The authors would like to thank the referees for constructive comments to improve the paper. This work has received funding from the National Fusion Programme of the Hellenic Republic--General Secretariat for Research and Innovation.